# Greedy Low-Rank Gradient Compression for Distributed Learning with Convergence Guarantees

### v2

@@ Appendix C Expanded example of non-contractive compressor
[…]
@@
[…]
However, since $\tau$ is typically on the order of hundreds, the term $2^{-2\tau}$ becomes astronomically small. For example, when $\tau>100$, $2^{-2\tau}<2^{-200}\approx 6.2\times 10^{-61},$ which is well below the smallest positive normalized value in IEEE-754 single-precision arithmetic (approximately $1.18\times 10^{-38}$). Consequently, $2^{-2\tau}$ underflows to zero in FP32, allowing us to assume $02^{-2\tau}=$ 0 and thus $\alpha=0$, in direct contradiction of the contraction property required by Assumption 1.
Input: $N$ nodes, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, initial weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and initial error buffer $\bm{E}_{-1}^{(i)}=\bm{0}\in\mathbb{R}^{\min(m,n)\times\max(m,n)}$ for node $i\in[N]$.
Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ Appendix D Detailed Implementations of GreedyLore
To make illustration of the algorithm simpler, we replace selection of left or right matrix multiplication with a transpose-based operation on gradients. After each local gradient $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})$ is computed immediately, each node transposes the gradient to $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})^{\top}\in\mathbb{R}^{n\times m}$ when $m>n$ and keep the shape $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})\in\mathbb{R}^{m\times n}$ when $m\leq n$ Then communication and compression are performed on this transposed form $n\times m$. Once the global compressed gradients are aggregated and reconstructed, they are transposed back to the original shape $m\times n$ for further update on optimizer state. Algorithm 4 summarizes the complete procedure.
Input: $N$ nodes, learning rate $\gamma$, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, $\beta_{1},\beta_{2}$ for Adam. Error buffer $\bm{E}_{-1}=\bm{0}\in\mathbb{R}^{m\times n}$, weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$, state for subspace optimizer $\bm{M}_{-1}={0},\bm{V}_{-1}={0}\in\mathbb{R}^{r\times n}$ and projection matrix $\bm{P}_{-1}\in\mathbb{R}^{m\times r}$ with $m\leq n$.
Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ Appendix F Missing Experimental Details
Pre-training tasks on CIFAR datasets We pre-train ResNet-18 models on CIFAR-10 and CIFAR-100 datasets for both 40 epochs on a 4 × 4090 24GB GPUs cluster. The experiments are configured with global batch size 4 × 32 in PyTorch DDP framework. We start the compression at 500 iterations for CIFAR-10 and 4 000 iterations for CIFAR-100 following the warm-up convention in PowerSGD[31]. We use max learning rate of $45e-3,5e-$ 4 with cosine annealing scheduler, subspace switching frequency $750,1200$ respectively for training on CIFAR-10 and CIFAR-100 datasets. For low-rank algorithm in 3, the compress rank is set to 64 in all settings.

### v4

@@ Appendix C Expanded example of non-contractive compressor
$\|\mathcal{C}(\nabla f(x_{\tau},y_{\tau}))\|_{F}^{2}=\bigl(L\,2^{-\tau}\bigr)^{2}=L^{2}\,2^{-2\tau},$
$\|\nabla f(x_{\tau},y_{\tau})\|_{F}^{2}=\bigl(L\,2^{-\tau}\bigr)^{2}+\bigl(\tfrac{L}{2}\bigr)^{2}=L^{2}\Bigl(2^{-2\tau}+\tfrac{1}{4}\Bigr).$
@@
$\alpha=\frac{\|\mathcal{C}(\nabla f(x_{\tau},y_{\tau}))\|_{F}^{2}}{\|\nabla f(x_{\tau},y_{\tau})\|_{F}^{2}}=\frac{L^{2}\,2^{-2\tau}}{L^{2}\bigl(2^{-2\tau}+\tfrac{1}{4}\bigr)}=\frac{2^{-2\tau}}{2^{-2\tau}+\tfrac{1}{4}}.$
However, since $\tau$ is typically on the order of hundreds, the term $2^{-2\tau}$ becomes astronomically small. For example, when $\tau>100$, $2^{-2\tau}<2^{-200}\approx 6.2\times 10^{-61},$ which is well below the smallest positive normalized value in IEEE-754 single-precision arithmetic (approximately $1.18\times 10^{-38}$). Consequently, $2^{-2\tau}$ underflows to zero in FP32, allowing us to assume $2^{-2\tau}=0$ and thus $\alpha=0$, in direct contradiction of the contraction property required by Assumption 1.
Input: NN nodes, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, initial weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$ and initial error buffer $\bm{E}_{-1}^{(i)}=\bm{0}\in\mathbb{R}^{\min(m,n)\times\max(m,n)}$ for node $i\in[N]$.
2pt Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ Appendix D Detailed Implementations of GreedyLore
To make illustration of the algorithm simpler, we replace selection of left or right matrix multiplication with a transpose-based operation on gradients. After each local gradient $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})$ is computed immediately, each node transposes the gradient to $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})^{\top}\in\mathbb{R}^{n\times m}$ when $m>n$ and keep the shape $\nabla F_{i}(\bm{X}_{t};{\bm{\xi}}_{t}^{(i)})\in\mathbb{R}^{m\times n}$ when $m\leq n$ . Then communication and compression are performed on this transposed form $n\times m$. Once the global compressed gradients are aggregated and reconstructed, they are transposed back to the original shape $m\times n$ for further update on optimizer state. Algorithm 4 summarizes the complete procedure.
Input: NN nodes, learning rate $\gamma$, number of total iterations $T$, subspace changing frequency $\tau$, rank $r$, $\beta_{1},\beta_{2}$ for Adam. Error buffer $\bm{E}_{-1}=\bm{0}\in\mathbb{R}^{m\times n}$, weight $\bm{X}_{0}\in\mathbb{R}^{m\times n}$, state for subspace optimizer $\bm{M}_{-1}={0},\bm{V}_{-1}={0}\in\mathbb{R}^{r\times n}$ and projection matrix $\bm{P}_{-1}\in\mathbb{R}^{m\times r}$ with $m\leq n$.
2pt Output: Sequence of model weights $\{\bm{X}_{t}\}_{t=0}^{T+1}$.
@@ Appendix F Missing Experimental Details
Pre-training tasks on CIFAR datasets We pre-train ResNet-18 models on CIFAR-10 and CIFAR-100 datasets for both 40 epochs on a 4 × 4090 24GB GPUs cluster. The experiments are configured with global batch size 4 × 32 in PyTorch DDP framework. We start the compression at 500 iterations for CIFAR-10 and 4 000 iterations for CIFAR-100 following the warm-up convention in PowerSGD[31]. We use max learning rate of 5e3,5e45e-3,5e-4 with cosine annealing scheduler, subspace switching frequency $750,1200$ respectively for training on CIFAR-10 and CIFAR-100 datasets. For low-rank algorithm in 3, the compress rank is set to 64 in all settings.